Use a generic D curve of the form P=a+bQ

At a P = $36 Q = 4

At a P = $27 Q = 7
Answer #1

P. 236 yey=4 P2 = 27, 20: P-Pi = 0-0, 80 Pc - Pg_De-R2 P-360 36-27 -4 4-7 P-36 = 9 P-36 -H -3 P-36 = -3 (0-4) 9-36 = 30 +12 P

y = -3x + 48.. R2 = 1 Axis Title + 0 Axis Title

I have provided answer with both method manual as well as Excel.

I hope my efforts will be fruitful to you...?